# HoneySat: A Network-based Satellite Honeypot Framework

### v2

@@ III Threat Model
Following Fig. 2, we assume that an adversary willing to compromise a satellite can only interact with the space segment simulation by sending TCs from the ground segment first. To gain initial access to the ground segment, an adversary needs to connect via one of the exposed network protocols discussed in Sec. II-E, namely, VNC, Telnet, TCP/IP, etc., which correspond to the operational protocols used in real satellite missions. From there, an adversary may try to launch different commands to take full control and/or compromise the services offered by the satellite’s mission as depicted in Fig. 3. Finally, in this paper, the modeling of physical radio communication between the space and ground segments, which is commonly used in practice, is considered out of scope and left for future work. The threat model is further referenced in Sec. IV-E.
@@ 1. DO-1
    Capability to Capture Rich Interaction Data. As explained in Sec. I, the number one objective of any honeypot is to capture interaction data from which we can derive knowledge on adversaries’ TTPs. As such, our first objective for HoneySat is to have the capability to capture rich interaction data.
@@ 3. DO-3
    Provide Extensibility and Customizability. A framework’s main purpose is to provide generic functionality that can be customized to meet the user’s needs. In HoneySat’s case, we must be able to support multiple SmallSats. For example, a particular SmallSat may use CSP or CCSDS ecosystems. Additionally, each SmallSat mission has a particular orbit that must be customizable. As such, our third objective is for HoneySat to provide extensibility and customizability.
@@ 2. DP-2
    Realistic Simulation. As we discussed in Sec. II-B, satellites keep track of their orbit location, the position of the sun, among other details. These details must be simulated by our honeypot; otherwise, they may alert adversaries that they are interacting with a fake system. For this reason, we selected realistic simulation as our second design principle. This design principle is based on design objective DO-2.
@@ 3. DP-3
    Modularity. As we explained in Sec. II, satellites are complex and diverse systems. There is no one-size-fits-all solution. To tackle this problem, we designed HoneySat to be modular. Modularity allows us to insert, remove, or change any part of our honeypot framework without implementing a new honeypot from scratch. This design principle is based on our objective DO-3.
We now describe HoneySat’s architecture design and how this architecture integrates the design principles described above. At the highest level, our framework’s architecture, depicted in Fig. 2, consists of two sets of simulations, the space segment simulations and the ground segment simulations. Table II illustrates how HoneySat’s simulation components match the components of a real satellite mission.
TABLE II: Correspondence between real satellite mission components and HoneySat’s design components.
Segment Real Component HoneySat Component
Ground Remote Operation Protocols VNC, Telnet
Ground Mission Control Software Modified MCS
Ground Ground Station Control Software Gpredict
Ground Web Interface Web Interface
Ground Ground Station Radio Simulator
Space Flight Software F.S. Runtime + Services
Space On Board Computer Docker Container
Space Platform Satellite Simulator
Space Payload Satellite Simulator
@@ IV-C Ground Segment Design
2) Ground Software. As discussed in Sec. II-C, the ground software includes mission control software (MCS) and ground station control software (GSCS). We leverage existing ground software, allowing HoneySat to provide a high-interaction simulation; thereby following design principle DP-1.
@@ IV-D Space Segment Design
2) Satellite Simulator. One of the biggest challenges when designing HoneySat was simulating all the physical processes, e.g., attitude, that satellites need to know. The Satellite Simulator is designed to solve this problem by simulating all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver, that a satellite requires to function. The purpose of the Satellite Simulator is to abstract away the various hardware peripherals that communicate with the flight software by simulating these peripherals’ data. For example, suppose that a particular flight software was implemented to utilize a specific lithium-ion battery pack. In that case, the Satellite Simulator’s power system simulation can simulate the basic behavior of the battery. We designed the Satellite Simulator to be modular so that new subsystems and sensors can be easily added, modified, or disabled according to our design principle DP-3.
@@ IV-E Theory of Operation
An adversary gets initial access via one of the Exposed Network Protocols ①. On interacting with the Simulated Ground Segment, an attacker is presented with access to the Ground Software ②. The configuration of this ground software is defined by the Ground Configuration, which allows the ground software to look like one of many different missions from its protocol ecosystem to the attacker. The attacker can then interact with the Ground Software, while their actions are reported to the Logging Repository. They might want to try to gain more privileges on the ground segment or try to send TC to the space segment. Instead of deploying a ground station with an RF transmitter and associated hardware, we employ the Radio Simulator ③ to handle all simulated radio frequency communications.
When an attacker uses the ground segment to send a valid TC or raw packets, the Radio Simulator checks whether the satellite configured in the Satellite Personality is currently passing over the designated ground station location. If so, the Radio Simulator forwards the TC to the Flight Software Runtime ④. When the Flight Software Runtime receives a command, it will run the corresponding Flight Software Service to compute a response ⑤. In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator ⑥. The Satellite Personality contains the configuration used by the Satellite Simulator. Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker ⑦-⑫, and is also recorded in our Logging Repository.
@@ V-A Satellite Simulator Implementation
We implemented the Satellite Simulator as a Python object-oriented programming application.
We implemented 11 sensors, 4 subsystems, 1 satellite state, and 1 interface, based on our 6 simulations. These modules are generic enough so that we do not have to heavily modified them to support other satellites. Overall, satellite architectures, subsystems, and sensors are varied, so modularity is necessary when constructing a simulation framework. Many of the basic simulations and corresponding sensors can be shared, especially in a honeypot where attackers lack detailed information that would enable them to distinguish between small differences in sensor values. Further details
can be found in Appendix A-B.
@@ V-B Generic CSP Mission Honeypot
Following our design, we implemented HoneySat to support the CSP ecosystem. This implementation provides a generic CSP-based mission honeypot that can be customized to simulate a specific CPS-based satellite mission.
@@ V-B1 Ground Segment Implementation
1) Exposed Network Protocols.
We implemented four exposed entry points. A VNC server, a Telnet server, a Web server, and ZeroMQ-based access to the CSP Network. We implemented the VNC server using TigerVNC [64] and used PyZMQ [65] for the Raw Packet Access. The Telnet server was implemented using Python to expose the command-line interface of the Ground Segment Software (GSS) on the network. The web interface allows for the customization of text presented to the attacker through configuration. Using the above entry points, adversaries can interact with the GSS and the space segment simulations.
@@
2) Ground Software. To implement the MCS software for our CSP-based honeypot, we require an MCS that supports the CSP ecosystem. We leveraged the SUCHAI mission control software (SUCHAI MCS). This software is a node on the CSP network that has a command line-based interface that allows basic MCS functionality like scheduling telecommands or saving downlinked telemetry. This is a good choice for the generic CSP MCS as it can be passed off as the CLI of commercial ground segment appliances like the GomSpace NanoCom MS100 [46]. Given HoneySat being a generic honeypot, we can set mission-specific parameters through the ground configuration. Additionally, we provide a desktop environment accessible via VNC, complete with supporting applications such as a configured Gpredict for the satellite mission. For certain deployments, this environment also includes a mission-specific GUI tool, both of which appear as open windows on the desktop.
3) Radio Simulator. The Radio Simulator was implemented as a ZMQ Hub [66]. The ZMQ Hub functions as a router for the CSP network simulating both the ground part of the CSP network including the ground station and the space segment. Depending on whether the satellite is currently passing, packets to the space segment are forwarded to the Flight Software Runtime.
We decided to use ZeroMQ as it is the protocol to use for transmitting CSP packets over TCP/IP [66].
4) Ground Configuration. We implemented the ground configuration as settings that can be modified through a Docker compose file explained in Appendix A-C.
5) Logging Repository. This implementation uses the MongoDB database [67] to aggregate and store logs. The web interface records login attempts, while the Telnet server logs inputs on a per-connection basis. Additionally, network traffic to and from the host running HoneySat is captured.
@@ V-B2 Space Segment Implementation
1) Flight Software Runtime. Although most of the flight software frameworks presented in Section II-D are open-source projects, it is difficult to find a real satellite flight software implementation that is available for study and modification [12]. Thus, we use the University of Chile’s SUCHAI Flight Software (SUCHAI FS) [52], which has been deployed in four satellite missions [68]. The SUCHAI FS provides flight software based on the CSP ecosystem that already features a mode where it can be run on a POSIX OS.
2) Flight Software Services. The SUCHAI FS uses libCSP [55] and thus features the libCSP default services that present a generic attack surface. Furthermore, there is a set of commands commonly used in satellites as a basic set of tasks to control common satellite operations such as communications. There are 96 of these commands and parameters that are parsed from CSP packets using a text-based protocol. We modified the command handlers to use the Satellite Simulator instead of commanding subsystems and querying real sensors.
3) Satellite Personality. The satellite personality was implemented as a Python class that holds multiple variables related to the space segment. The current implementation includes almost 50 configurable variables, such as the satellite’s name, the battery’s nominal cell capacity, etc.
@@
Each logged data entry is created with a customized tag and a timestamp. The collected data is structured JSON [69] format and stored in the MongoDB database.
Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A-C.
@@ 3. Q-3
    Can HoneySat be extended to simulate different SmallSat missions and support hardware-in-the-loop simulations?
@@ VI-B TTP Interaction Experiment
This experiment seeks to answer Q-1 by quantifying the interactions provided by HoneySat. To achieve this, we leveraged the SPACE-SHIELD matrix (Version 2.0) [63] discussed in Sec. II-F. SPACE-SHIELD provides a collection of adversary tactics and techniques for space systems. In this experiment, we determined the number of tactics and techniques that HoneySat supports. SPACE-SHIELD consists of 14 tactics and 62 techniques. However, not all of them are applicable to a virtual, network-based honeypot such as HoneySat. For example, the technique Compromise Hardware Supply Chain involves “replacing a hardware component in the supply chain with a custom or counterfeit part” which is out of the scope of a virtual honeypot. Taking this into consideration, 14 tactics are applicable to HoneySat. From these tactics, HoneySat can feasibly offer up to 38 techniques as possible interactions for adversaries.
@@
Experiment Methodology. We designed one interaction or exploit for each of the 38 applicable techniques for our honeypot. The interactions involved ground segment simulations such as the web interface. The exploits were simple, using one TC, or complex with multiple TCs involved.
For example, the technique System Service Discovery (ID T1007)’s description reads “Adversaries may try to gather information about registered local system services. Adversaries may obtain information about services using tools and OS utility commands. Adversaries may use the information from System Service Discovery during automated discovery to shape follow-on behaviors, including whether or not the adversary fully infects the target and/or attempts specific actions.”
Based on this description, we designed the following exploit to capture information about the satellite’s running processes:
@@
Experiment Results. We successfully crafted and ran an interaction or exploit on HoneySat for 33 techniques out of the feasible 38 as depicted in Table III. The remaining 5 techniques were not implemented due to limitations in HoneySat’s implementation. For example, the Retrieve TT&C master/session keys (T2015.002) technique depends on a cryptographic protocol implementation which HoneySat’s flight software currently does not support. This is not surprising, as satellite flight software rarely if ever does [12].
Due to space limitations, we do not describe all the exploits and interactions here. However, the complete exploit and interaction list is available in Table VIII in Appendix A.
@@
TABLE III: Tactics and techniques supported by HoneySat compared to the ones included in the SPACE-SHIELD matrix that are applicable to virtual honeypots.
@@ VI-C SmallSat Operators Survey
Evaluating the realism of a satellite honeypot is challenging for two main reasons. First, as discussed in Sec. I, satellites, including SmallSats, are very diverse. Second, there is no established metric or tool, such as Nmap’s OS detection [70], to quantify the level of realism of our honeypot.
@@
TABLE IV: Summary of questions in Section VI-C, satellite honeypot operation tasks of the survey.
@@ VI-C1 Survey Structure
We divided our survey into five sections. The first section collected the participants’ background information including demographic data. The second section focused on their professional experience, including whether they worked in industry, government, or academia and to self-report their satellite-related and cybersecurity skills (e.g., satellite hardware engineering). The third section probed participants’ satellite operation experience, including how many missions they had operated and which tools they used.
The fourth survey section involved giving participants hands-on access to a live instance of HoneySat. In this section, participants were asked to perform 4 hands-on tasks that correspond to specific satellite mission operations using a live instance of HoneySat as depicted in Fig. 4. Each hands-on task was designed to feature different components of HoneySat by replicating a real-world satellite mission operation, as discussed in Section II. After participants completed each hands-on task, they answered questions designed to assess how realistic they found different aspects of the honeypot. Table IV lists the satellite operations included in the survey.
Finally, in section five, once participants had interacted with multiple elements of our honeypot via these tasks, the survey concluded with participants answering questions about HoneySat’s overall realism and deception. Questions related to both the honeypot operation tasks and the overall evaluation used a 5-point Likert scale [71] to measure the participants’ reactions.

### v3

@@ III Threat Model
Following Fig. 2, we assume that an adversary willing to compromise a satellite can only interact with the space segment simulation by sending TCs from the ground segment first. To gain initial access to the ground segment, an adversary needs to connect via one of the exposed network protocols discussed in Sec. II-E, which correspond to the operational protocols used in real satellite missions. From there, an adversary may try to launch different commands to take full control and/or compromise the services offered by the satellite’s mission as depicted in Fig. 4. Finally, in this paper, the modeling of physical radio communication between the space and ground segments, which is commonly used in practice, is considered out of scope and left for future work. The threat model is further referenced in Sec. IV-E.
@@ 1. DO-1
    Capability to Capture Rich Interaction Data. As explained in Sec. I, the number one objective of any honeypot is to capture interaction data from which we derive knowledge on adversaries’ TTPs. As such, HoneySat’s first objective is to capture rich interaction data.
@@ 3. DO-3
    Provide Extensibility and Customizability. A framework’s main purpose is to provide generic functionality that can be customized to meet the user’s needs. In HoneySat’s case, we must be able to support multiple SmallSats. For example, a particular SmallSat may use CSP or CCSDS ecosystems. As such, HoneySat’s third objective is extensibility and customizability.
@@ 2. DP-2
    Realistic Simulation. As discussed in Sec. II-B, satellites track their orbit location, among others. These details must be simulated; otherwise, they may alert adversaries that they are interacting with a fake system. This design principle is based on design objective DO-2.
@@ 3. DP-3
    Modularity. As we explained in Sec. II, satellites are complex and diverse systems. To tackle this problem, we designed HoneySat to be modular. This design principle is based on our objective DO-3.
We now describe HoneySat’s architecture design and how this architecture integrates the design principles described above. At the highest level, our framework’s architecture, depicted in Fig. 2, consists of two sets of simulations, the space segment simulations and the ground segment simulations. Table VII illustrates how HoneySat’s simulation components match the components of a real satellite mission.
@@ IV-C Ground Segment Design
2) Ground Software. As discussed in Sec. II-C, the ground software includes mission control software (MCS) and ground station control software (GSCS). We leverage existing ground software, e.g., Gpredict, allowing HoneySat to provide a high-interaction simulation following design principle DP-1.
@@ IV-D Space Segment Design
Figure 3: High-level architecture of the Satellite Simulator. Due to space limitations, we do not show all the available sensors and simulations.
2) Satellite Simulator. One of the biggest challenges when designing HoneySat was simulating all the physical processes, e.g., attitude, that satellites need to know. The Satellite Simulator solves this problem by simulating all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver.
As shown in Fig. 3, the Satellite Simulator includes five module types: simulation modules (orange), sensor modules (blue), subsystem state modules (green), satellite state modules (gray) and interface modules (yellow).
Simulation Modules. These modules are abstractions of real-world processes required for a realistic satellite simulation (DP-2), e.g., the orbital simulation which uses orbital mechanics to calculate data like the satellite’s orbital position.
These simulation modules can communicate with each other to facilitate proper functionality. For example, the power simulation queries the orbital simulation for data to determine if the satellite is positioned properly to draw power from its solar panels. The Satellite Simulator includes six simulations, the orbital, rotation, power system, thermal, magnetic, and payload simulations.
  • Orbital Simulation. The orbital simulation calculates the satellite’s orbit [vallado2012two]. It uses the TLE data configured in the satellite personality and it generates multiple values such as, latitude, and longitude of the satellite. The payload, EPS, and magnetic simulations rely on data from this simulation.
  • Rotation Simulation. This simulation calculates the satellite’s attitude changes to provide its orientation by using a reference frame fixed to the satellite body and a non-rotational reference frame. The relation between the two reference frames is calculated using the conservation of angular momentum and the rigid-body Euler equation [euler]. This simulation’s data is used by payloads like the Red, Green, Blue (RGB) camera to point towards Earth’s surface.
  • Power System Simulation. This simulation manages the satellite’s power collection, consumption, and distribution. It tracks the battery capacity, the power draw and simulates battery charging whenever the orbital simulation tells it that the satellite is exposed to sunlight.
  • Thermal Simulation. This simulation calculates the satellite’s temperature based on the total thermal energy and a user-defined specific heat capacity and uses a radiation loss formula [stefan-boltzmann] to calculate the thermal radiation emission.
  • Magnetic Simulation. This simulation tracks and analyzes the interactions between the Earth’s magnetic field and the satellite’s own magnetic environment. It communicates with the orbital and rotation simulations to determine the satellite’s position and orientation to calculate the Earth’s magnetic field components.
  • RGB Camera Simulation. This simulation replicates the functionality of an RGB camera pointed at Earth. It uses Earth observation satellite imagery from the U.S. Geological Survey (USGS) [geological2022earthexplorer]. If a capture image command is issued, it will select and return an image. This simulation communicates with the orbital simulation to determine if satellite is in the presence of the sun.
@@ IV-E Theory of Operation
An adversary gets initial access via one of the Exposed Network Protocols 1. On interacting with the Simulated Ground Segment, an attacker is presented with access to the Ground Software 2. The configuration of this ground software is defined by the Ground Configuration, which allows the ground software to look like one of many different missions from its protocol ecosystem to the attacker. The attacker can then interact with the Ground Software, while their actions are reported to the Logging Repository. They might want to try to gain more privileges on the ground segment or try to send TC to the space segment. Instead of deploying a ground station with an RF transmitter and associated hardware, we employ the Radio Simulator 3 to handle all simulated radio frequency communications.
When an attacker uses the ground segment to send a valid TC or raw packets, the Radio Simulator checks whether the satellite configured in the Satellite Personality is currently passing over the designated ground station location. If so, the Radio Simulator forwards the TC to the Flight Software Runtime 4. When the Flight Software Runtime receives a command, it will run the corresponding Flight Software Service to compute a response 5. In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator 6. The Satellite Personality contains the configuration used by the Satellite Simulator. Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker 7- 12, and is also recorded in our Logging Repository.
@@ V-A Satellite Simulator Implementation
We implemented the Satellite Simulator as a Python object-oriented programming application. We implemented 11 sensors, 4 subsystems, 1 satellite state, and 1 interface, based on our 6 simulations.
@@ V-B Generic CSP Mission Honeypot
We implemented HoneySat to support the CSP ecosystem providing a generic CSP-based honeypot that can simulate any specific CSP-based mission.
@@ V-B1 Ground Segment Implementation
1) Exposed Network Protocols. We implemented a VNC server, a Telnet server, a Web server, and ZeroMQ-based access to the CSP Network. We implemented the VNC server using TigerVNC [ossman2024tigervnc] and used PyZMQ [granger2024pyzmq] for the raw packet access. The Telnet server was implemented using Python to expose the command-line interface of the Ground Segment Software. The web interface allows for the customization of text presented to the attacker through configuration.
@@
2) Ground Software. To implement the MCS we leveraged the SUCHAI mission control software (SUCHAI MCS). This software is a node on the CSP network that has a command line-based interface that allows basic MCS functionality like scheduling telecommands or saving downlinked telemetry. Additionally, we provide a desktop environment accessible via VNC that includes Gpredict and the SUCHAI MCS. Finally, we implemented a radio link and network delay simulation using a sleep function whenever the MCS receives data from the space segment simulations.
3) Radio Simulator. The Radio Simulator was implemented as a ZMQ Hub [ProtocolStackCubesatZMQ]. The ZMQ Hub functions as a router for the CSP network simulating both the ground part of the CSP network including the ground station and the space segment. Depending on whether the satellite is currently passing, packets to the space segment are forwarded to the Flight Software Runtime. Additionally, the Radio Simulator dynamically simulates packet loss depending on the predicted elevation angle of the satellite (provided by the Satellite Simulator’s orbital simulation). For example, if an attacker sends a telecommand, the Radio Simulator will use a probability value based on the elevation angle and decide to drop or forward the packet.
4) Ground Configuration. We implemented the ground configuration as settings that can be modified through a Docker compose file explained in Appendix A.
5) Logging Repository. This implementation uses the MongoDB database [mongodb2024mongodb] to aggregate and store logs. The web interface records login attempts, while the Telnet server logs inputs on a per-connection basis. Additionally, network traffic to and from the host running HoneySat is captured. The logs are secured at points that are not observable by attackers, e.g., different Docker container, shown in Fig. 6, and the isolated MongoDB database is accessed by an append-only user.
@@ V-B2 Space Segment Implementation
1) Flight Software Runtime. We use the SUCHAI Flight Software (SUCHAI FS) [gonzalez2019architecture], which has been deployed in four satellite missions [Garrido2023] and provides flight software based on the CSP ecosystem.
2) Flight Software Services. The SUCHAI FS uses libCSP [LibcspLibcsp2024] and thus features the libCSP default services that present a generic attack surface. SUCHAI FS includes a set of telecommands which we modified to use the Satellite Simulator instead of querying real sensors.
3) Satellite Personality. The satellite personality was implemented as a Python class that holds multiple variables related to the space segment. It includes almost 50 configurable variables, such as the satellite’s name.
@@
Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A.
V-C Summary of Existing and Newly Developed Software
HoneySat’s implementation involved the integration of existing, modified, and newly developed software. As-is software includes TigerVNC, MongoDB, and Gpredict. Modified software includes the SUCHAI FS, SUCHAI MCS, the Telnet server, and the ZeroMQ Hub. Newly developed software includes the Satellite Simulator, including all 6 simulations, the web interface, the flight software services for FS integration, the satellite personality, the ground configuration, and all the necessary framework infrastructure, including Docker files, integration scripts, and configuration files.
@@ 3. Q-3
    Can HoneySat be extended to support different SmallSat missions and hardware-in-the-loop simulations?
@@ VI-B TTP Interaction Experiment
This experiment seeks to answer Q-1 by quantifying the interactions provided by HoneySat. To achieve this, we leveraged the SPACE-SHIELD matrix (Version 2.0) [esa2024space-shield] which provides a collection of adversary tactics and techniques for space systems. We determined the number of tactics and techniques that HoneySat supports. SPACE-SHIELD consists of 14 tactics and 62 techniques. However, not all of them are applicable to a virtual, network-based honeypot such as HoneySat. For example, the technique Compromise Hardware Supply Chain involves “replacing a hardware component in the supply chain with a custom or counterfeit part” which is out of the scope of a virtual honeypot. Taking this into consideration, 14 tactics are applicable to HoneySat. From these tactics, HoneySat can feasibly offer up to 38 techniques as interactions for adversaries.
@@
Experiment Methodology. We designed one interaction or exploit for each of the 38 applicable techniques for our honeypot. The interactions involved ground segment simulations such as the web interface. The exploits were simple, using one TC, or complex with multiple TCs involved. For example, the technique System Service Discovery (ID T1007) involves adversaries obtaining information about services using tools and OS utility commands. Based on this description, we designed the following exploit to capture information about the satellite’s running processes:
@@
Experiment Results. We successfully crafted and ran an interaction or exploit on HoneySat for 33 techniques out of the feasible 38 as depicted in Table II. The remaining 5 techniques were not implemented due to limitations in HoneySat’s implementation. For example, the Retrieve TT&C master/session keys (T2015.002) technique depends on a cryptographic protocol implementation which HoneySat’s flight software currently does not support. Due to space limitations, we do not describe all the exploits and interactions here. However, the complete exploit and interaction list is available in Table VIII in Appendix B.
@@
TABLE II: Tactics and techniques supported by HoneySat.
@@ VI-C SmallSat Operators Survey
Evaluating the realism of a satellite honeypot is challenging for two main reasons. First, as discussed in Sec. I, satellites, including SmallSats, are very diverse. Second, there is no established metric or tool, such as Nmap’s OS detection [nmap2025nmap], to quantify the level of realism of our honeypot.
@@
TABLE III: Summary of questions in Section VI-C, satellite honeypot operation tasks of the survey.
@@ VI-C1 Survey Structure
We divided our survey into five sections. The first section collected the participants’ background information, e.g., demographic data, the second section focused on their professional experience, the third section probed participants’ satellite operation experience, e.g., how many missions they had operated.
The fourth survey section involved giving participants hands-on access to a live instance of HoneySat. In this section, participants were asked to perform 4 hands-on tasks that correspond to specific satellite mission operations using a live instance of HoneySat as depicted in Fig. 5. Each hands-on task was designed to feature different components of HoneySat by replicating a real-world satellite mission operation, as discussed in Section II. After participants completed each hands-on task, they answered questions designed to assess how realistic they found different aspects of the honeypot. Table III lists the satellite operations included in the survey.
Finally, in section five, once participants had interacted with multiple elements of our honeypot, the survey concluded with participants answering questions about HoneySat’s overall realism and deception. Questions related to both the honeypot operation tasks and the overall evaluation used a 5-point Likert scale [likert2025qualtrics] to measure the participants’ reactions.